OpenStudy (anonymous):

How do you find the Pythagorean triple and the give is 3,4, and x.

OpenStudy (rulnick):

x=5

OpenStudy (kinggeorge):

simply solve the equation \[\sqrt{3^2+4^2} =x\]

OpenStudy (rulnick):

pythag triples satisfy x^2 + y^2 = z^2 with x, y, and z all integers

OpenStudy (anonymous):

3^2+4^2 = 9 + 16 = 25 = 5^2 3-4-5 is a Pythagorean triple!

OpenStudy (anonymous):

You guys are so incredible